# Elastic Plateau–Rayleigh instability in soft cylinders: Surface elasticity and periodic beading

### v1

@@ 1 Introduction
Similarly, experiments show that elastic filaments can undergo a surface-tension-driven instability, resulting in a periodic sequence of bulges along the elastic body (Matsuo and Tanaka,, 1992; Zuo et al.,, 2005; Naraghi et al.,, 2007; Mora et al.,, 2010), a phenomenon known as beading. This elastic analogue of the Plateau–Rayleigh instability has been investigated in recent years; see, for instance, (Mora et al.,, 2010; Ciarletta and Ben Amar,, 2012; Taffetani and Ciarletta, 2015b, ; Xuan and Biggins,, 2016). However, existing theoretical models predict only the formation of an isolated bulge (Lestringant and Audoly,, 2020; Fu et al.,, 2021), which contradicts experimental observations, leaving the phenomenon unexplained.
In most of these papers, surface tension in solids is treated as a constant quantity, similarly to fluids (Mora et al.,, 2010). Although this might be an acceptable approximation in many cases, studies have shown that surface tension in solids has an elastic nature, i.e. its value depends on the deformation of the surface, see fig. 1. The pioneering idea behind this phenomenon is due to Shuttleworth, (1950), who postulated a linear dependence of the surface stress with respect to the surface deformation.
@@ 2.1 Notation and basic kinematics
where $\left[\mathsf{\Sigma}^{+}\right]^{-1}$ is the pseudoinverse of $\mathsf{\Sigma}$ obtained substituting each non-zero entry on the diagonal of $\mathsf{\Sigma}$ with its reciprocal value. By construction, we observe that (Yu and Fu,, 2025; Javili et al.,, 2014)
@@ 2.2 Pre-stretched elastic surfaces: balance equations
Figure 2: Representation of the multiplicative decomposition $\mathsf{F}_{s}=\mathsf{F}^{e}_{s}\mathsf{F}^{p}_{s}$.
@@
where $\nabla_{s}\cdot$ denotes the surface divergence operator (see Appendix A for a definition). We remark that eq. 8 provides the boundary condition for the classical balance equation of linear momentum of a three dimensional body
@@
In isothermal conditions, if we assume the existence of a surface strain energy density $\psi_{s}=\psi_{s}(\mathsf{F}_{s})$, standard thermodynamic considerations allow us to write (Dehghany et al.,, 2020)
@@
This equation characterizes hyperelastic material surfaces, similarly to classical hyperelastic materials. Using the multiplicative decomposition of the deformation gradient eq. 7, we assume that the surface strain energy density depends on $\mathsf{F}_{s}^{e}$ only, i.e. there exits a function $\psi_{s0}$ such that
$\psi_{s}(\mathsf{F}_{s})=\psi_{s0}\left(\mathsf{F}_{s}{\left(\mathsf{F}_{s}^{p}\right)}^{-1}\right).$
@@
$\mathsf{P}_{s}=\frac{\partial\psi_{s}}{\partial\mathsf{F}_{s}^{e}}\left(\mathsf{F}_{s}^{p}\right)^{-T}.$ (11)
@@ 3 Solid elastic cylinder surrounded by a pre-tensioned elastic surface
Let $\mathscr{B}_{0}\subset\mathbb{R}^{3}$ be the reference configuration representing the cylinder, with $R_{0}$ denoting its radius. We introduce the referential cylindical coordinates $(R,\,\Theta,\,Z)$ and the corresponding vector basis $(\boldsymbol{E}_{R},\,\boldsymbol{E}_{\Theta},\,\boldsymbol{E}_{Z})$. We assume that its axial length is much greater than the radius, so that we can assume $\mathscr{B}_{0}$ to be infinite in the direction $\boldsymbol{E}_{Z}$.
@@
$\psi_{s}=\frac{\mu_{s}}{2}\bigg{(}I_{s}^{e}-2-2\ln J_{e}\bigg{)}+\frac{\Lambda_{s}}{2}\left(\frac{1}{2}\left(J_{e}^{2}-1\right)-\ln J_{e}\right),$ (16)
where $\mu_{s}$ is the surface shear modulus and $\Lambda_{s}$ modulates surface extensibility. The quantity $I_{s}^{e}$ is defined as $I_{s}^{e}=\operatorname{tr}\left(\left(\mathsf{F}_{s}^{e}\right)^{T}\mathsf{F}_{s}^{e}\right)$, while $J_{e}=\operatorname{det}_{s}\mathsf{F}_{s}^{e}$ represents the elastic part of the surface Jacobian $J_{s}$.
@@ 4 Stability analysis
In this section we conduct a linear stability analysis of the cylindrical configurations. First, we show that the cylinder always admits a solution with cylindrical symmetry.
@@ 4.1 Cylindrical solution
Finally, we can find the expression of the pressure field $p$ by enforcing the boundary condition eq. 12b. Using eqs. 21 and 22 we get
@@ 4.2 Incremental relations
where we assume summation over repeated indices. The incremental counterpart of eqs. 12a and 12b and eq. 13 are given by
@@ 4.3 Axisymmetric solutions of the incremental problem and linear stability analysis
It is convenient to proceed with a non-dimensionalization, identifying the dimensionless groups that control the bifurcation. In particular, we choose $R_{0}$ and $\mu_{s}$ as characteristic length and stiffness of the system, respectively. This choice allows us to identify the following dimensionless quantities
$\widehat{\mu}=\frac{\mu R_{0}}{\mu_{s}},$ $\widehat{\Lambda}_{s}=\frac{\Lambda_{s}}{\mu_{s}},$ $\widehat{k}=kR_{0}.$
We observe that $\mu_{s}/\mu$ can be interpreted as the counterpart of the classical elasto-capillary length of the system (Style et al.,, 2017), while $\widehat{\Lambda}_{s}$ measures the extensibility of the free surface. Equation (36) can now be rewritten in dimensionless form as
$\widehat{\varphi}\left(\widehat{k},\widehat{\mu},\lambda,\widehat{\Lambda}_{s},\lambda_{p}\right)=0,$ (37)
where $\widehat{\varphi}$ is the nondimensionalized counterpart of the function $\varphi$.
@@ 4.4 Results of the linear stability analysis
Figure 5: Plot of $\widehat{\mu}_{\text{cr}}$ (turquoise) and $\widehat{k}_{\text{cr}}$ (blue) against $\lambda_{p}$ for $\lambda=1$ (fig. 5a) and $\lambda\neq 1$ (fig. 5b). Here, $\widehat{\Lambda}_{s}=40$.
In this section, we report and discuss the results of the linear stability analysis. We refer to the first mode that becomes unstable as the critical buckling mode. The corresponding dimensionless critical wavenumber is denoted by $\widehat{k}_{\text{cr}}$.
We first explore the stability of the cylindrical configuration with respect to the control parameter $\widehat{\mu}$, analysing the effect of surface pre-stretch. In all the studied cases, as we decrease $\widehat{\mu}$, the critical mode has a non-zero wavenumber, see fig. 4. The marginal stability curves in fig. 4 show a similar trend as we change the parameters. Specifically, we observe a decrease in the critical thresholds for $\widehat{\mu}$ as $\lambda_{p}$ approaches $1$. Moreover, axial stretching appears to stabilize the cylinder: a bifurcation occurs at larger values of $\widehat{\mu}$ when $\lambda=1.4$ with respect to the case with $\lambda=1$, as shown in fig. 4. In fig. 5 we report the trend of the critical wavenumber and of the critical shear modulus. Our results show that $\widehat{\mu}_{\text{cr}}$ monotonically decreases as $\lambda_{p}$ increases.
We now analyse the effect of surface extensibility. The results are reported in fig. 6. As the surface becomes less extensible, i.e. as $\widehat{\Lambda}_{s}$ grows, also $\widehat{\mu}_{\text{cr}}$ monotonically grows. From fig. 7 we can notice that this trend is linear for both $\lambda=1$ and $\lambda=1.4$. In particular, the axial strain has a stabilizing effect on the cylinder, increasing $\widehat{\mu}_{\text{cr}}$ as the axial stretch grows. Furthermore, we observe that the critical wavenumber rapidly increases for small $\widehat{\Lambda}_{s}$, and saturates to a constant value when the surface is nearly inextensible. From a physical standpoint, these results suggest that the periodic beading pattern can be triggered more easily when the surface is stiffer and nearly inextensible. We also remark that, as before, the critical wavenumber predicted by the linear stability analysis is finite and non-zero, in agreement with the beading pattern observed in the experiments (Mora et al.,, 2010).
@@
Figure 7: Plot of $\widehat{\mu}_{\text{cr}}$ (turquoise) and $\widehat{k}_{\text{cr}}$ (blue) against $\widehat{\Lambda}_{s}$ for $\lambda=1$ (fig. 7a) and $\lambda\neq 1$ (fig. 7b). Here, $\lambda_{p}=0.8$.
(a) $\lambda=1$.
(b) $\lambda=1.4$.
[…]
(a) $\lambda=1$.
(b) $\lambda=1.4$.
Figure 9: Plot of $\lambda_{p}^{cr}$ (turquoise) and $\widehat{k}_{\text{cr}}$ (blue) against $\widehat{\mu}$ for $\lambda=1$ (fig. 9a) and $\lambda\neq 1$ (fig. 9b). Here, $\widehat{\Lambda}_{s}=40$.
Figure 10: Plot of the axial stretch $\lambda$ versus the dimensionless wavenumber $k$. The arrow denotes the direction of growth of $\lambda_{p}$ with uniform steps of amplitude 0.05, from 0.5 to 0.7.
Figure 11: Plot of $\lambda_{p}^{cr}$ (turquoise) and $\widehat{k}_{\text{cr}}$ (blue) against $\lambda$. Here, we set $\widehat{\Lambda}_{s}=10$ and $\widehat{\mu}=0.8$.
@@ 5 Numerical simulations
In this section, we detail the numerical methods and present the results of the simulations for the nonlinear boundary value problem (12a)-(18).
@@ 5.1 Weak formulation and finite-element approximation
while homogeneous Neumann conditions are assumed for the remaining component of the traction.
@@
where $\boldsymbol{y}$ represents the state of a (physical) system and $\eta$ is the control parameter of the bifurcation problem. In our application, $\boldsymbol{y}=\left(\bm{u},\,p\right)$ and $\eta$ is one of the dimensionless parameters of eq. 37.
Continuation algorithms allow to find the region of the plane $\left(\boldsymbol{y},\eta\right)$ where eq. 38 is satisfied. Specifically, we exploited a pseudo-arclength continuation algorithm, which requires to adopt the control parameter $\eta$ as an additional unknown of the system. As a consequence, we must add a further equation to the problem. Assuming that we know that $\boldsymbol{f}(\boldsymbol{y}_{j},\,\eta_{j})=0$, we look for a couple $(\boldsymbol{y}_{j+1},\,\eta_{j+1})$ that satisfies eq. 38 and a constraint of the form
@@
$\mathcal{V}_{0}$ =$=\Big{\{}\bm{v}\in\left[H^{1}\left(\Sigma\right)\right]^{2}:v_{Z}=0\text{ on }\Gamma_{1}\cup\Gamma_{2}\cup\Gamma_{3}\Big{\}},$

### v2

@@ 1 Introduction
Similarly, experiments show that elastic filaments can undergo a surface-tension-driven instability, resulting in a periodic sequence of bulges along the elastic body (Matsuo and Tanaka,, 1992; Zuo et al.,, 2005; Naraghi et al.,, 2007; Mora et al.,, 2010), a phenomenon known as beading. This elastic analogue of the Plateau–Rayleigh instability has been investigated in recent years, see, for instance, Mora et al., (2010); Ciarletta and Ben Amar, (2012); Taffetani and Ciarletta, 2015b ; Xuan and Biggins, (2016), but existing theoretical models predict only the formation of an isolated bulge (Lestringant and Audoly,, 2020; Fu et al.,, 2021).
Several studies have suggested that the experimentally observed finite wavelength may result from the sensitivity of the system to imperfections; see, for example, Taffetani and Ciarletta, 2015a ; Giudici and Biggins, (2020). However, the experiments by Matsuo and Tanaka (Matsuo and Tanaka,, 1992) demonstrate that the wavelength consistently scales with the radius of the cylinder. This scaling behavior indicates that wavelength selection is a robust and intrinsic feature of the instability, rather than a consequence of random surface imperfections. The occurrence of finite wavelength selection appears to be a common feature across several experiments, as reviewed in the introduction of Fu et al., (2021).
In most of the previously cited papers, surface tension in solids is treated as a constant quantity, similarly to fluids (Mora et al.,, 2010). Although this might be an acceptable approximation in many cases, studies have shown that surface tension in solids has an elastic nature, i.e. its value depends on the deformation of the surface, see fig. 1. The pioneering idea behind this phenomenon is due to Shuttleworth, (1950), who postulated a linear dependence of the surface stress with respect to the surface deformation.
@@ 2.1 Notation and basic kinematics
where $\left[\mathsf{\Sigma}^{+}\right]^{-1}$ is the pseudoinverse of $\mathsf{\Sigma}$ obtained by substituting each non-zero entry on the diagonal of $\mathsf{\Sigma}$ with its reciprocal value. By construction, we observe that (Yu and Fu,, 2025; Javili et al.,, 2014)
@@ 2.2 Pre-stretched elastic surfaces: balance equations
Figure 2: Representation of the multiplicative decomposition of the surface defomation gradient tensor $\mathsf{F}_{s}=\mathsf{F}^{e}_{s}\mathsf{F}^{p}_{s}$.
@@
where $\mathsf{P}$ is the (bulk) Piola-Kirchhoff stress tensor, while $\nabla_{s}\cdot$ denotes the surface divergence operator (see Appendix A for a definition). We remark that eq. 8 provides the boundary condition for the classical balance equation of linear momentum of a three dimensional body
@@
In isothermal conditions, if we assume the existence of a surface strain energy density per unit reference area $\psi_{s}=\psi_{s}(\mathsf{F}_{s})$, standard thermodynamic considerations allow us to write (Dehghany et al.,, 2020)
@@
This equation characterizes hyperelastic material surfaces, similarly to classical hyperelastic materials. In order to construct the strain energy density of the material in the presence of a pre-stretch, a careful treatment is required.
The boundary layer generated by capillary forces is very small, so we are allowed to consider only stretching effects and neglect bending energy, as in the theoretical framework of Gurtin and Murdoch, (1975). It is important to track the variations of the thickness of the material surface in the process. We assume that the material surface represents the limit of a thin layer of a three-dimensional incompressible material. Let $H$ and $H_{r}$ be the (infinitesimal) thicknesses of the material surface in the reference and in the relaxed state, respectively. We recall that $\operatorname{det}_{s}\mathsf{F}_{s}^{p}$ describes the change in surface area from the reference to the relaxed state, so that
$\mathrm{d}S_{r}=\operatorname{det}_{s}\mathsf{F}_{s}^{p}\,\mathrm{d}S,$ (11)
where $\mathrm{d}S_{r}$ and $\mathrm{d}S$ are the infinitesimal area element in the relaxed and reference state, respectively. Thus, from the incompressibility of the material we have
$H=H_{r}\operatorname{det}_{s}\mathsf{F}_{s}^{p}.$
We consider the strain energy density $\psi_{s}^{r}$ per unit surface area in the relaxed state. We introduce $W_{s}^{r}$, defined as
$W_{s}^{r}(\mathsf{F}_{s}^{e})\coloneqq\frac{\psi_{s}^{r}(\mathsf{F}_{s}^{e})}{H_{r}}.$
Here, $W_{s}^{r}$ can be regarded as strain energy density per unit volume averaged across the material surface thickness. By eq. 11, introducing $\psi_{s0}\left(\mathsf{F}_{s}^{e}\right)=W_{s}^{r}\left(\mathsf{F}_{s}^{e}\right)H$, we get
$\psi_{s}^{r}\left(\mathsf{F}_{s}{\left(\mathsf{F}_{s}^{p}\right)}^{-1}\right)\,\mathrm{d}S_{r}=\psi_{s0}\left(\mathsf{F}_{s}{\left(\mathsf{F}_{s}^{p}\right)}^{-1}\right)\,\mathrm{d}S=\psi_{s}(\mathsf{F}_{s})\,\mathrm{d}S.$
@@
$\mathsf{P}_{s}=\frac{\partial\psi_{s0}}{\partial\mathsf{F}_{s}^{e}}\left(\mathsf{F}_{s}^{p}\right)^{-T}.$ (12)
@@ 3 Solid elastic cylinder surrounded by a pre-tensioned elastic surface
Let $\mathscr{B}_{0}\subset\mathbb{R}^{3}$ be the reference configuration representing the cylinder, with $R_{0}$ denoting its radius. We introduce the referential cylindrical coordinates $(R,\,\Theta,\,Z)$ and the corresponding vector basis $(\boldsymbol{E}_{R},\,\boldsymbol{E}_{\Theta},\,\boldsymbol{E}_{Z})$. We assume that its axial length is much greater than the radius, so that we can assume $\mathscr{B}_{0}$ to be infinite in the direction $\boldsymbol{E}_{Z}$.
@@
$\psi_{s0}(\mathsf{F}_{s}^{e})=\frac{\mu_{s}}{2}\bigg{(}I_{s}^{e}-2-2\ln J_{e}\bigg{)}+\frac{\Lambda_{s}}{2}\left(\frac{1}{2}\left(J_{e}^{2}-1\right)-\ln J_{e}\right),$ (17)
where $\mu_{s}$ is the surface shear modulus and $\Lambda_{s}$ modulates surface extensibility (the higher $\Lambda_{s}$, the more inextensible). The quantity $I_{s}^{e}$ is defined as $I_{s}^{e}=\operatorname{tr}\left(\left(\mathsf{F}_{s}^{e}\right)^{T}\mathsf{F}_{s}^{e}\right)$, while $J_{e}=\operatorname{det}_{s}\mathsf{F}_{s}^{e}$ represents the elastic part of the surface Jacobian $J_{s}$.
@@ 4 Stability analysis
In this section we conduct a linear stability analysis of the cylindrical configurations with respect to axisymmetric perturbations with a wavenumber $k$. In what follows, we show that the body always admits a cylindrical solution, and we perform a stability analysis. Readers mainly interested in results are suggested to go directly to section 4.4.
@@ 4.1 Cylindrical solution
Finally, we can find the expression of the pressure field $p$ by enforcing the boundary condition eq. 13b. Using eqs. 22 and 23 we get
@@ 4.2 Incremental relations
where we assume summation over repeated indices. The incremental counterpart of eqs. 13a and 13b and eq. 14 are given by
@@ 4.3 Axisymmetric solutions of the incremental problem and linear stability analysis
In the following section we perform a non-dimensionalization of eq. 37 and present and discuss the results of the linear stability analysis.
@@ 4.4 Results of the linear stability analysis
(a) $\lambda=1$.
(b) $\lambda=1.4$.
[…]
First, we identify the non-dimensional quantities that govern the behaviour of the system. In particular, we choose $R_{0}$ and $\mu_{s}$ as characteristic length and stiffness of the system, respectively. This choice allows us to identify the following dimensionless quantities
$\widehat{\mu}=\frac{\mu R_{0}}{\mu_{s}},$ $\widehat{\Lambda}_{s}=\frac{\Lambda_{s}}{\mu_{s}},$ $\widehat{k}=kR_{0},$
representing the dimensionless shear modulus, the dimensionless surface extensibility, and the dimensionless wavenumber, respectively. We observe that $\mu_{s}/\mu$ can be interpreted as the counterpart of the classical elasto-capillary length of the system (Style et al.,, 2017), while $\widehat{\Lambda}_{s}$ measures the extensibility of the free surface.
(a) $\lambda=1$.
(b) $\lambda=1.4$.
Figure 5: Plot of the critical bifurcation thresholds of the dimensionless shear modulus $\widehat{\mu}_{\text{cr}}$ (turquoise) and of the dimensionless wavenumber $\widehat{k}_{\text{cr}}$ (blue) against the pre-stretch parameter $\lambda_{p}$ for unitary (fig. 5a) and nonunitary (fig. 5b) axial strain, respectively. Here, the dimensionless surface extensbility $\widehat{\Lambda}_{s}$ is set to 40.
@@
Figure 7: Plot of the critical bifurcation thresholds of the dimensionless shear modulus $\widehat{\mu}_{\text{cr}}$ (turquoise) and of the dimensionless wavenumber $\widehat{k}_{\text{cr}}$ (blue) against the dimensionless surface extensibility $\widehat{\Lambda}_{s}$ for unitary (fig. 7a) and nonunitary (fig. 7b) axial strain, respectively. Here, the pre-stretch parameter $\lambda_{p}$ is set to 0.8.
In the following, we refer to the first mode that becomes unstable as the critical buckling mode. The corresponding dimensionless critical wavenumber is denoted by $\widehat{k}_{\text{cr}}$.
We first explore the stability of the cylindrical configuration with respect to control parameter $\widehat{\mu}$, analysing the effect of surface pre-stretch. In all the studied cases, as we decrease $\widehat{\mu}$, the critical mode has a non-zero wavenumber, see fig. 4. The marginal stability curves in fig. 4 show a similar trend as we change the parameters. Specifically, we observe a decrease in the critical thresholds for $\widehat{\mu}$ as $\lambda_{p}$ approaches $1$. Moreover, axial stretching appears to stabilize the cylinder: a bifurcation occurs at larger values of $\widehat{\mu}$ when $\lambda=1.4$ with respect to the case with $\lambda=1$, as shown in fig. 4. In fig. 5 we report the trend of the critical wavenumber and of the critical shear modulus. Our results show that $\widehat{\mu}_{\text{cr}}$ monotonically decreases as $\lambda_{p}$ increases.
We now analyse the effect of surface extensibility. The results are reported in fig. 6. As the surface becomes less extensible, i.e. as $\widehat{\Lambda}_{s}$ grows, also $\widehat{\mu}_{\text{cr}}$ monotonically grows. From fig. 7 we can notice that this trend is linear for both $\lambda=1$ and $\lambda=1.4$. In particular, the axial strain has a stabilizing effect on the cylinder, increasing $\widehat{\mu}_{\text{cr}}$ as the axial stretch grows. Furthermore, we observe that the critical wavenumber rapidly increases for small $\widehat{\Lambda}_{s}$, and saturates to a constant value when the surface is nearly inextensible. From a physical standpoint, these results suggest that the periodic beading pattern can be triggered more easily when the surface is stiffer and nearly inextensible. We also remark that, as before, the critical wavenumber predicted by the linear stability analysis is finite while the system remains stable for longer wavelengths ($\widehat{k}\rightarrow 0$). Thus, periodic patterning appears to be favourable compared to isolated bulging, in agreement with the observations in the experiments (Mora et al.,, 2010).
@@
Figure 9: Plot of the critical bifurcation thresholds of the pre-stretch parameter $\lambda_{p}^{cr}$ (turquoise) and of the dimensionless wavenumber $\widehat{k}_{\text{cr}}$ (blue) against the dimensionless shear modulus $\widehat{\mu}$ for unitary (fig. 9a) and nonunitary (fig. 9b) axial strain, respectively. Here, the dimensionless surface extensibility $\widehat{\Lambda}_{s}$ is set to 40.
Figure 10: Plot of the axial stretch $\lambda$ versus the dimensionless wavenumber $k$. The arrow denotes the direction of growth of the pre-stretch parameter $\lambda_{p}$ with uniform steps of amplitude 0.05, from 0.5 to 0.7.
Figure 11: Plot of the critical bifurcation thresholds of the pre-stretch parameter $\lambda_{p}^{cr}$ (turquoise) and of the dimensionless wavenumber $\widehat{k}_{\text{cr}}$ (blue) against the axial strain $\lambda$. Here, we set the dimensionless surface extensibility $\widehat{\Lambda}_{s}$ to 10 and the dimensionless shear modulus $\widehat{\mu}$ to 0.8.
@@ 5 Numerical simulations
In this section, we detail the numerical methods and present the results of the simulations for the nonlinear boundary value problem (13a)-(19), both close to the bifurcation threshold and in the post-buckling regime. Readers mainly interested in results are suggested to go directly to section 5.2.
@@ 5.1 Weak formulation and finite-element approximation
while homogeneous Neumann conditions are assumed for the remaining component of the traction. In particular, on $\Gamma_{1}$ and $\Gamma_{2}$ roller-type boundary conditions are imposed: the axial displacement is prescribed to enforce the mean stretch $\lambda$, while the radial displacement is left free under zero traction. These conditions effectively act as symmetry boundaries, enabling the modelling of a segment of an infinite, uniformly stretched cylinder.
@@
where $\boldsymbol{y}$ represents the state of a (physical) system and $\eta$ is the control parameter of the bifurcation problem. In our application, $\boldsymbol{y}=\left(\bm{u},\,p\right)$ and $\eta$ is one of the dimensionless parameters $\widehat{\mu},\lambda$ and $\lambda_{p}$.
Continuation algorithms allow us to find the region of the plane $\left(\boldsymbol{y},\eta\right)$ where eq. 38 is satisfied. Specifically, we exploited a pseudo-arclength continuation algorithm, which requires the adoption of the control parameter $\eta$ as an additional unknown of the system. As a consequence, we must add a further equation to the problem. Assuming that we know that $\boldsymbol{f}(\boldsymbol{y}_{j},\,\eta_{j})=0$, we look for a couple $(\boldsymbol{y}_{j+1},\,\eta_{j+1})$ that satisfies eq. 38 and a constraint of the form
@@
$\mathcal{V}_{0}$ =$=\Big{\{}\bm{v}\in\left[H^{1}\left(\Sigma\right)\right]^{2}:v_{R}=0\text{ on }\Gamma_{2}\cup\Gamma_{3},\,v_{Z}=0\text{ on }\Gamma_{1}\Big{\}},$